To decrease production costs, a company suggests replacing one of its manufacturing equipment with a newer, more efficient model. This four year project will result in reduced manufacturing costs which, in turn, would allow a reduction in the price of its finished product. The current equipment can be sold today for $1,000,000 net. A brand-new equipment retails almost $3,250,000; however, it can be purchased for $3,000,000. Funding for this purchase will include proceeds from the sale of the old equipment. In addition, accounts payable are expected to increase by $1,500,000 today, and fully reverse in year 4.

The new equipment will be in operation beginning in year two. As the old equipment will be offline in year 1, the company forecasts lost revenues of $550,000 in year 1. The cost savings in years 2, 3 and 4 are estimated at $600,000, $950,000, and $1,000,000, respectively. The company's cost of capital and tax-rate remain unchanged. What is the initial outlay for this project, the project’s FCF for years 1 through 4, and the project’s NPV? The company is concerned that the estimated cost of capital is too high. It adjusts it's Beta and computes a new cost of capital of 5%. Using this, what is this project’s NPV?

Exhibits

MACRS 5 Schedule

Year 1: 20%

Year 2: 32%

Year 3: 19.2%

Year 4: 11.52%

Year 5: 11.52%

Cost of Capital

WACC =

Cost of Equity = Risk-Free Rate + β *(Equity Risk Premium)

After-Tax Cost of Debt = (1-Tax Rate) * Pre-Tax Cost of Debt

Risk-Free Rate (10-Year U.S. Treasury) = 3%

The Equity Risk Premium = 4.5%

Tax Rate: 40%

Company's beta (β) = 1.2

Company's Market Value of Equity / Total Capital ratio = 100%

Company's Market Value of Debt / Total Capital = 0%
Answer #1

Answer:

Initial outlay of the project = Cost of new equipment - Sale proceeds net of old equipment - Increase in accounts payables

= 3000000 - 1000000 - 1500000

= $500,000

Initial outlay of the project = $500,000

Year 1 to Year 4 Cash flows

Year 1 cash flow = Lost revenue * (1 - Tax rate) = - 550000 * (1 - Tax rate) = -550000 * (1 - 40%) = - $330,000

Year 2 cash flow = cost savings * (1 - Tax rate) + Depreciation tax shield = 600000 * (1 - 40%) + 3000000 * 20% * 40%

= $600,000

Year 3 cash flow = 950000 * (1 - 40%) + 3000000 * 32% * 40% = $954,000

Year 4 cash flow = cost savings * (1 - Tax rate) + Depreciation tax shield - reinvestment of working capital

= 1000000 * (1 - 40%) + 3000000 * 19.2% * 40% - 1500000

= - $669,600

NPV calculation at 5% cost of capital:

NPV = -330000 / (1 + 5%) + 600000 / (1 + 5%) 2 + 954000 / (1 + 5%) 3 - 669600 / (1 + 5%) 4 - 500000

= $3,151.46

Using 5% cost of capital project’s NPV = $3,151.46

Year 0 1 2 3 4 Cost of ($3,000,000) Sale proceeds of old machine (net) $1,000,000 $1,500,000 equipment new Working capital ($